Calculate the angular momentum of the following systems. Refer to Worksheet 19 for the

moment of inertia of highly-symmetric objects. For all items, you need to convert from

rpm to rad/s first.


Bicycle wheel with mass 2.0 kg and diameter 0.63 m rotating at a rate of 2.00

rpm. Assume that the bicycle wheel is a thin-walled hollow cylinder.


1
Expert's answer
2021-02-15T00:57:43-0500

The moment of inertia of the thin-walled hollow cylinder:



"I = mr^2"


where "m = 2kg" is the mass, and "r = 0.63m\/2 = 0.315m" is the radius of the wheel.


"I = 2\cdot (0.315)^2 =0.19845 \space kg\cdot m^2"


The angular frequency:


"\omega = 2\pi\dfrac{2}{60} \approx 0.21\space rad/s"

The angular momentum:



"L = I\omega\\
L \approx 0.042\space kg\cdot m^2/s"

Answer. 0.042 kg*m^2/s.